# ASVspoof 5: Design, Collection and Validation of Resources for Spoofing, Deepfake, and Adversarial Attack Detection Using Crowdsourced Speech

### v2

@@ 7.1 Baselines, MOS estimator, and metrics
The SASV systems are an ASV-CM fusion-based system [119] (B03, middle panel of Figure 3), and an end-to-end system (B4 [120], right panel of Figure 3). B03 is the SASV 2022 challenge baseline [119, 121] and is a LLR-based fusion [122] of the ASV and AASIST CM baselines described above. B04, based on MFA-Conformer [123], extracts a single embedding from the input waveform and produces a single output score. It is trained in three stages: pre-training for speaker classification; copy synthesis training [124] with adapted SASV loss functions; in-domain fine-tuning using ASVspoof 5 training data.
@@
The CM EER is estimated using a set of bona fide and spoofed trial scores. SASV performance estimates are expressed in terms of the architecture-agnostic detection cost function (a-DCF) [126, 33], a cost-based performance metric designed specifically for the SASV task. While a binary classification task, there are now three classes, namely trials involving bona fide targets (positive class) and then bona fide non-targets and spoofed target trials (negative class). Lower a-DCF values indicate better performance. [^21] Implementations of all baseline systems and evaluation metrics are accessible from the ASVspoof 5 Github repository. Further details for the evaluation metrics are available in the ASVspoof 5 challenge evaluation plan [33].
@@ 7.2 Impacts upon ASV
Attacks in the evaluation set result in some of the highest EERs. Of the 16 attacks, 14 lead to EERs higher than 20%. A19, also a legacy, unit-selection based TTS system, is once again the attack which provokes the highest EER. These results are in line with previously reported findings [8, 35, 108], a challenge recognized since the very first ASVspoof challenge [2, 129]. A decade after, this persistent challenge is worthwhile emphasising: the most effective attack algorithms are not necessarily those implemented with the latest technology. Legacy TTS systems, based on the concatenation of bona fide speech segments extracted from the utterances of a target speaker are particularly effective.
Pre-trained TTS attacks A17, A28, and A29 give the next highest EERs, all above 31%. [^22] A small portion of A17’s training data consists of utterances from 12 target speakers but does not overlapped with the ASVspoof 5 evaluation set (Section 3.4). The ASV EER computed on the bona fide and spoofed data of the ‘seen’ 12 target speakers is 42.51%, which is lower than that of the ‘unseen‘ (44.65%). The high ASV EER provoked by A17 is not due to the ‘seen’ target speakers. A21 and A22, both trained using only data specified in the ASVspoof 5 contributors’ protocol, also provoke high EERs. VC-based attacks A24, A25 and A26 are among the least effective. The effect of encoding/compression for the evaluation set is illustrated to the right in Figure 9, which shows score distributions for bona fide target and spoofed/deepfake trials. As one might expect, once again, encoding/compression reduces the differences between the distributions, provoking higher EERs.
@@ Appendix B Dendrogram of ASVspoof 5 data
where ||$\cdot$|| ⋅ | denotes the number of observations and $c(\boldsymbol{a},\boldsymbol{b})$ is cosine distance of vectors $\boldsymbol{a}$ and $\boldsymbol{b}$. The similarity matrix $\boldsymbol{S}$ and is used for hierarchical agglomerative clustering, using the Ward’s minimum variance criterion as the clustering objective. The figure is rendered using the Seanborn clustermap API [109].
@@ Appendix C DET Curves
Figure 10: DET curves of ASV and CM baselines illustrating performance across different attacks in the evaluation set. Solid lines are zero-effort (from non-target speakers) and non-adversarial attacks. Dashed lines are adversarial attacks, including Malafide, Malacopula, and their combination. The color of the dashed line is consistent with the non-adversarial attack counterpart. EER operating points of non-adversarial, Malafide, Malacopula, and combination of Malafide and Malacopula attacks are marked with ,,$\bullet$,$\blacktriangle$,$\times$ , ▲ , ×, and $\blacksquare$, respectively. Numbers listed in the legend are EER values. The order of attacks in the legend is sorted based on the EER.

### v3

@@ 7.1 Baselines, MOS estimator, and metrics
The SASV systems are an ASV-CM fusion-based system [119] (B03, middle panel of Figure 3), and an end-to-end system (B4 [120], right panel of Figure 3). B03 is the SASV 2022 challenge baseline [119, 121] and is an LLR-based fusion [122] of the ASV and AASIST CM baselines described above. B04, based on MFA-Conformer [123], extracts a single embedding from the input waveform and produces a single output score. It is trained in three stages: pre-training for speaker classification; copy synthesis training [124] with adapted SASV loss functions; in-domain fine-tuning using ASVspoof 5 training data.
@@
The CM EER is estimated using a set of bona fide and spoofed trial scores. SASV performance estimates are expressed in terms of the architecture-agnostic detection cost function (a-DCF) [126, 34], a cost-based performance metric designed specifically for the SASV task. While the SASV makes a binary decision, there are now three classes, namely trials involving bona fide targets (positive class) and then bona fide non-targets and spoofed target trials (negative class). Lower a-DCF values indicate better performance. [^21] Implementations of all baseline systems and evaluation metrics are accessible from the ASVspoof 5 Github repository. Further details for the evaluation metrics are available in the ASVspoof 5 challenge evaluation plan [34].
@@ 7.2 Impacts upon ASV
Attacks in the evaluation set result in some of the highest EERs. Of the 16 attacks, 14 lead to EERs higher than 20%. A19, also a legacy, unit-selection based TTS system, is once again the attack which provokes the highest EER. These results are in line with previously reported findings [9, 36], a challenge recognized since the very first ASVspoof challenge [2, 3]. A decade after, this persistent challenge is worthwhile emphasising: the most effective attack algorithms are not necessarily those implemented with the latest technology. Legacy TTS systems, based on the concatenation of bona fide speech segments extracted from the utterances of a target speaker are particularly effective.
Pre-trained TTS attacks A17, A28, and A29 give the next highest EERs, all above 31%. [^22] A small portion of A17’s training data consists of utterances from 12 target speakers but does not overlapped with the ASVspoof 5 evaluation set (Section 3.4). The ASV EER computed on the bona fide and spoofed data of the ‘seen’ 12 target speakers is 42.51%, which is lower than that of the ‘unseen’ (44.65%). The high ASV EER provoked by A17 is not due to the ‘seen’ target speakers. A21 and A22, both trained using only data specified in the ASVspoof 5 contributors’ protocol, also provoke high EERs. VC-based attacks A24, A25 and A26 are among the least effective. The effect of encoding/compression for the evaluation set is illustrated to the right in Figure 9, which shows score distributions for bona fide target and spoofed/deepfake trials. As one might expect, once again, encoding/compression reduces the differences between the distributions, provoking higher EERs.
@@ Appendix B Dendrogram of ASVspoof 5 data
where ||$\cdot$|| ⋅ | denotes the number of observations and $c(\boldsymbol{a},\boldsymbol{b})$ is cosine distance of vectors $\boldsymbol{a}$ and $\boldsymbol{b}$. The similarity matrix $\boldsymbol{S}$ is used for hierarchical agglomerative clustering, using the Ward’s minimum variance criterion as the clustering objective. The figure is rendered using the Seanborn clustermap API [109].
@@ Appendix C DET Curves
Figure 10: DET curves of ASV and CM baselines illustrating performance across different attacks in the evaluation set. Solid lines are zero-effort (from non-target speakers) and non-adversarial attacks. Dashed lines are adversarial attacks, including Malafide, Malacopula, and their combination. The colour of the dashed line is consistent with the non-adversarial attack counterpart. EER operating points of non-adversarial, Malafide, Malacopula, and combination of Malafide and Malacopula attacks are marked with ,,$\bullet$,$\blacktriangle$,$\times$ , ▲ , ×, and $\blacksquare$, respectively. Numbers listed in the legend are EER values. The order of attacks in the legend is sorted based on the EER.